OpenStudy (anonymous):

A bag contains tiles with the letters P-R-O-B-A-B-I-L-I-T-Y. Tyler chooses a tile without looking and doesn’t replace it. He chooses a second tile without looking. What is the probability that he will choose the letter I both times?

OpenStudy (anonymous):

\[\frac{2}{11}\times \frac{1}{10}\]

OpenStudy (anonymous):

a)1/55 b)2/55 c) 2/121 d)3/121

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thank u